# Adaptive Pareto-Optimal Token Merging for Edge Transformer Models in Semantic Communication

### v1

@@ I Introduction
The rapid evolution of wireless networks is fundamentally transforming how data and intelligence are delivered, processed, and experienced at the network edge. As we look towards sixth-generation (6G) systems, the emphasis is shifting from sheer connectivity and raw data throughput towards enabling intelligent, context-aware services that adapt in real time to users’ needs and environmental dynamics [1]. In this new paradigm, the concept of semantic communication, extracting and transmitting informative features that are directly relevant to downstream tasks, has emerged as a key enabler for bandwidth and energy-efficient edge intelligence [2].
Much of the existing work on semantic communication focuses on optimizing the delivery of feature representations or compressed signals, usually through compact convolutional neural networks or task-specific autoencoders to reduce the amount of data transmitted across the network [3, 4]. However, recent breakthroughs in deep learning, particularly with the rise of transformer architectures [5], have demonstrated remarkable advances in extracting semantically rich representations from high-dimensional multi-modal data such as images, audio, and text. Despite their promise, transformers remain computationally and communicatively intensive, which makes their practical deployment at the wireless edge challenging due to strict constraints on computation, memory, and uplink bandwidth [6].
@@
In this paper, we propose a training-free, multi-objective optimization framework for transformer-based semantic communication at the wireless edge. Our approach enables an edge device to process raw data using a pretrained transformer model, dynamically merging and reducing semantic tokens as they propagate through the model’s layers. The resulting compact set of tokens is then transmitted over a joint source-channel coded (JSCC) wireless link to a server for downstream task execution. We formulate the token merging process as a multi-objective optimization problem, jointly maximizing task performance and minimizing inference time. By employing Bayesian optimization [8] and constructing a Pareto front of optimal merging configurations, our method empowers users or network controllers to flexibly select the most suitable operating point for any given scenario, trading off accuracy and efficiency in real time, and doing so entirely without retraining.
@@ II Related Work
Qiao et al. introduced Token Communications (TokCom) [9], which defines semantic-level units based on transformer tokens, and uses pretrained models to generate and transmit context-aware tokens. Their approach shows improvements compared to previous work, however, improving the efficiency of the transformer model is not explored. Similarly, Xie et al. introduced large model-empowered semantic communication systems based on pretrained transformers [10], their results show improvements compared to previous baselines, however they highlight the need for further research to improve the deployment and usage of these computationally demanding models. Devoto et al. proposed an adaptive, trainable token pruning mechanism within a transformer and JSCC pipeline [11]. Their work requires retraining and is focused on dropping tokens entirely instead of merging them.
Bolya et al. introduced Token Merging (ToMe) [12], a training-free mechanism that merges similar transformer tokens during inference using a matching algorithm. Unlike pruning, ToMe reduces token redundancy by combining rather than dropping, achieving up to 2× inference speedup with minimal accuracy loss on Vision Transformers (ViTs). ToMe merges tokens based on similarity but uses a fixed merging schedule per layer and utilizes averaging to combine the tokens. Liu et al. proposed Adaptive Sparse ViT (AS-ViT) [13], that introduces learnable thresholds within self-attention layers to prune unimportant tokens dynamically during inference. Their scheme, trained with a budget-aware loss, can adapt pruning per sample and achieve significant speedups with minor accuracy degradation. However, this method requires retraining and focuses on token pruning instead of merging. Bayesian optimization for multi-objective latency-accuracy trade-offs has also been investigated in the context of neural architecture search [14], but this line of work has not been explored for the dynamic selection or merging of transformer tokens.
@@ IV Problem Statement
Transformer-based models have set the state-of-the-art across a wide range of tasks, but their high computational complexity presents a major obstacle for edge deployment. In particular, the self-attention mechanism in each transformer layer operates with a complexity of $\mathcal{O}(N^{2}d)$ [15]. As a result, inference with large $N$ is prohibitively expensive for resource-constrained edge devices. Therefore, if $N$ can be reduced during inference, significant savings in computation can be achieved without retraining.
@@ V-B Multi-Objective Bayesian Optimization for Pareto-Optimal Merging
Because the mappings $A(\boldsymbol{p})$ and $F(\boldsymbol{p})$ are a black-box, non-convex, high-dimensional, and costly to evaluate, we adopt Gaussian process (GP) Bayesian optimization to efficiently guide the search for Pareto-optimal configurations [16, 8]. In this approach, each objective is modeled as a realization of a GP prior:
@@
where $\mu(\boldsymbol{p})$ is the prior mean function and $k(\boldsymbol{p},\boldsymbol{p}^{\prime})$ is the covariance kernel. We choose the mean function to be constant and employ a Matérn-5/2 kernel with Automatic Relevance Determination (ARD) [17], defined in its simplified form as
@@
Here, $\sigma_{f}^{2}$ denotes the signal variance, which governs the overall variability of the function, while the length-scale parameters $\{\ell_{i}\}$ reflect the sensitivity of each merging proportion dimension to variations in the objective. Smaller length-scales indicate that small changes in $p_{i}$ can lead to rapid changes in the output, whereas larger length-scales imply smoother, less sensitive dependence. The ARD formulation thus enables the model to automatically learn which layers have the greatest impact on performance and computational cost. All kernel hyperparameters, including $\sigma_{f}^{2}$ and $\{\ell_{i}\}$, are estimated by maximizing the marginal likelihood of the observed data. To improve robustness and mitigate overfitting, Gamma priors are applied to the kernel scale and the observation noise variance, and additional priors are placed over the inverse squared length-scales [18].
Candidate configurations are selected by maximizing an acquisition function computed from the GP posterior predictive distributions. In the multi-objective setting, we adopt the logarithm of the Expected Hypervolume Improvement (EHVI) criterion [19], which quantifies the expected gain in the hypervolume dominated by the Pareto front when evaluating a candidate configuration. This criterion balances exploration of uncertain regions with exploitation of promising areas, enabling efficient approximation of the Pareto-optimal trade-offs. After a fixed budget of evaluations, the set of non-dominated configurations among those actually sampled provides an empirical approximation of the Pareto front as follows
$\widehat{\mathcal{F}}=\bigl{\{}\boldsymbol{p}\in\mathcal{S}_{T}\,\big{|}\,\nexists\,\boldsymbol{p}^{\prime}\in\ \mathcal{D}_{T}:A(\boldsymbol{p}^{\prime})\geq A(\boldsymbol{p})\wedge F(\boldsymbol{p}^{\prime})\leq F(\boldsymbol{p})\bigr{\}},$ (17)
@@ VI-A Experimental Setup
We evaluate our framework on the ImageNet-1k validation set using a ViT-Base/16 model pretrained with masked autoencoding [20, 21]. All transformer weights are frozen during inference. For each transformer layer, a merging proportion $p_{\ell}\in[0,0.3]$ specifies the fraction of tokens to merge. Gaussian process Bayesian optimization with a Matérn-5/2 kernel and ARD is conducted for 500 trials to identify Pareto-optimal configurations. The optimization is performed on a randomly sampled subset of 1000 validation images to reduce computational cost, while final evaluations are conducted on the full validation set. Each trial records top-1 classification accuracy and total FLOPs. The resulting configurations are then evaluated under both noiseless inference and noisy channel conditions. Specifically, an AWGN channel is simulated across SNR levels from -10 to 25 dB. We employ a lightweight convolutional JSCC encoder-decoder architecture similar to [3], trained separately on ImageNet embeddings. Baselines include the uncompressed model with no merging, uniform merging corresponding to the ToMe procedure applied with fixed proportions between 10% and 30%, and random merging schedules sampled uniformly. We report accuracy versus FLOPs trade-offs in the noiseless setting, end-to-end accuracy as a function of SNR, and throughput measured in images per second on an NVIDIA RTX 4080 GPU. Visual examples are also provided to illustrate the qualitative impact of token merging.  [^1] The source code, models and results will be made available at https://github.com/OmarErak/adaptive-token-merging-semcom
@@ VI-B Experimental Evaluation
VII Conclusions and Future Direction

### v2

@@ Abstract
footnotetext: An expanded version of this article can be found in[erak2025adaptive].
@@ I Introduction
The rapid evolution of wireless networks is fundamentally transforming how data and intelligence are delivered, processed, and experienced at the network edge. As we look towards sixth-generation (6G) systems, the emphasis is shifting from sheer connectivity and raw data throughput towards enabling intelligent, context-aware services that adapt in real time to users’ needs and environmental dynamics [6GVision]. In this new paradigm, the concept of semantic communication, extracting and transmitting informative features that are directly relevant to downstream tasks, has emerged as a key enabler for bandwidth and energy-efficient edge intelligence [chaccour2024less].
Much of the existing work on semantic communication focuses on optimizing the delivery of feature representations or compressed signals, usually through compact convolutional neural networks or task-specific autoencoders to reduce the amount of data transmitted across the network [bourtsoulatze2019deep, erak2024contrastive]. However, recent breakthroughs in deep learning, particularly with the rise of transformer architectures [vaswani2017attention], have demonstrated remarkable advances in extracting semantically rich representations from high-dimensional multi-modal data such as images, audio, and text. Despite their promise, transformers remain computationally and communicatively intensive, which makes their practical deployment at the wireless edge challenging due to strict constraints on computation, memory, and uplink bandwidth [letaief2021edge].
@@
In this paper, we propose a training-free, multi-objective optimization framework for transformer-based semantic communication at the wireless edge. Our approach enables an edge device to process raw data using a pretrained transformer model, dynamically merging and reducing semantic tokens as they propagate through the model’s layers. The resulting compact set of tokens is then transmitted over a joint source-channel coded (JSCC) wireless link to a server for downstream task execution. We formulate the token merging process as a multi-objective optimization problem, jointly maximizing task performance and minimizing inference time. By employing Bayesian optimization [frazier2018tutorialbayesianoptimization] and constructing a Pareto front of optimal merging configurations, our method empowers users or network controllers to flexibly select the most suitable operating point for any given scenario, trading off accuracy and efficiency in real time, and doing so entirely without retraining.
@@ II Related Work
Qiao et al. introduced Token Communications (TokCom) [qiao2025tokencommunicationsunifiedframework], which defines semantic-level units based on transformer tokens, and uses pretrained models to generate and transmit context-aware tokens. Their approach shows improvements compared to previous work, however, improving the efficiency of the transformer model is not explored. Similarly, Xie et al. introduced large model-empowered semantic communication systems based on pretrained transformers [10599117], their results show improvements compared to previous baselines, however they highlight the need for further research to improve the deployment and usage of these computationally demanding models. Devoto et al. proposed an adaptive, trainable token pruning mechanism within a transformer and JSCC pipeline [devoto2024adaptivesemantictokenselection]. Their work requires retraining and is focused on dropping tokens entirely instead of merging them.
Bolya et al. introduced Token Merging (ToMe) [bolya2023token], a training-free mechanism that merges similar transformer tokens during inference using a matching algorithm. Unlike pruning, ToMe reduces token redundancy by combining rather than dropping, achieving up to 2× inference speedup with minimal accuracy loss on Vision Transformers (ViTs). ToMe merges tokens based on similarity but uses a fixed merging schedule per layer and utilizes averaging to combine the tokens. Liu et al. proposed Adaptive Sparse ViT (AS-ViT) [10.24963/ijcai.2023/136], that introduces learnable thresholds within self-attention layers to prune unimportant tokens dynamically during inference. Their scheme, trained with a budget-aware loss, can adapt pruning per sample and achieve significant speedups with minor accuracy degradation. However, this method requires retraining and focuses on token pruning instead of merging. Bayesian optimization for multi-objective latency-accuracy trade-offs has also been investigated in the context of neural architecture search [eriksson2021latencyaware], but this line of work has not been explored for the dynamic selection or merging of transformer tokens.
@@ IV Problem Statement
Transformer-based models have set the state-of-the-art across a wide range of tasks, but their high computational complexity presents a major obstacle for edge deployment. In particular, the self-attention mechanism in each transformer layer operates with a complexity of $\mathcal{O}(N^{2}d)$ [keles2023computational]. As a result, inference with large $N$ is prohibitively expensive for resource-constrained edge devices. Therefore, if $N$ can be reduced during inference, significant savings in computation can be achieved without retraining.
@@ V-B Multi-Objective Bayesian Optimization for Pareto-Optimal Merging
Because the mappings $A(\boldsymbol{p})$ and $F(\boldsymbol{p})$ are a black-box, non-convex, high-dimensional, and costly to evaluate, we adopt Gaussian process (GP) Bayesian optimization to efficiently guide the search for Pareto-optimal configurations [rasmussen2003gaussian, frazier2018tutorialbayesianoptimization]. In this approach, each objective is modeled as a realization of a GP prior:
@@
where $\mu(\boldsymbol{p})$ is the prior mean function and $k(\boldsymbol{p},\boldsymbol{p}^{\prime})$ is the covariance kernel. We choose the mean function to be constant and employ a Matérn-5/2 kernel with Automatic Relevance Determination (ARD) [genton2001classes], defined in its simplified form as
@@
Here, $\sigma_{f}^{2}$ denotes the signal variance, which governs the overall variability of the function, while the length-scale parameters $\{\ell_{i}\}$ reflect the sensitivity of each merging proportion dimension to variations in the objective. Smaller length-scales indicate that small changes in $p_{i}$ can lead to rapid changes in the output, whereas larger length-scales imply smoother, less sensitive dependence. The ARD formulation thus enables the model to automatically learn which layers have the greatest impact on performance and computational cost. All kernel hyperparameters, including $\sigma_{f}^{2}$ and $\{\ell_{i}\}$, are estimated by maximizing the marginal likelihood of the observed data. To improve robustness and mitigate overfitting, Gamma priors are applied to the kernel scale and the observation noise variance, and additional priors are placed over the inverse squared length-scales [akiba2019optunanextgenerationhyperparameteroptimization].
Candidate configurations are selected by maximizing an acquisition function computed from the GP posterior predictive distributions. In the multi-objective setting, we adopt the logarithm of the Expected Hypervolume Improvement (EHVI) criterion [daulton2020differentiable], which quantifies the expected gain in the hypervolume dominated by the Pareto front when evaluating a candidate configuration. This criterion balances exploration of uncertain regions with exploitation of promising areas, enabling efficient approximation of the Pareto-optimal trade-offs. After a fixed budget of evaluations, the set of non-dominated configurations among those actually sampled provides an empirical approximation of the Pareto front as follows
$\widehat{\mathcal{F}}=\bigl\{\boldsymbol{p}\in\mathcal{S}_{T}\,\big|\,\nexists\,\boldsymbol{p}^{\prime}\in\ \mathcal{S}_{T}:A(\boldsymbol{p}^{\prime})\geq A(\boldsymbol{p})\wedge F(\boldsymbol{p}^{\prime})\leq F(\boldsymbol{p})\bigr\},$ (17)
@@ VI-A Experimental Setup
We evaluate our framework on the ImageNet-1k validation set using a ViT-Base/16 model pretrained with masked autoencoding [5206848, dosovitskiy2020image]. All transformer weights are frozen during inference. For each transformer layer, a merging proportion $p_{\ell}\in[0,0.3]$ specifies the fraction of tokens to merge. Gaussian process Bayesian optimization with a Matérn-5/2 kernel and ARD is conducted for 500 trials to identify Pareto-optimal configurations. The optimization is performed on a randomly sampled subset of 1000 validation images to reduce computational cost, while final evaluations are conducted on the full validation set. Each trial records top-1 classification accuracy and total FLOPs. The resulting configurations are then evaluated under both noiseless inference and noisy channel conditions. Specifically, an AWGN channel is simulated across SNR levels from -10 to 25 dB. We employ a lightweight convolutional JSCC encoder-decoder architecture similar to [bourtsoulatze2019deep], trained separately on ImageNet embeddings. Baselines include the uncompressed model with no merging, uniform merging corresponding to the ToMe procedure applied with fixed proportions between 10% and 30%, and random merging schedules sampled uniformly. We report accuracy versus FLOPs trade-offs in the noiseless setting, end-to-end accuracy as a function of SNR, and throughput measured in images per second on an NVIDIA RTX 4080 GPU. Visual examples are also provided to illustrate the qualitative impact of token merging.  [^1] The source code, models and results will be made available at https://github.com/OmarErak/adaptive-token-merging-semcom
@@ VI-B Experimental Evaluation
VII Conclusions and Future Directions